OpenStudy (anonymous):

Let f(x) = ln(1 + (1/x)). If g(x) = e^(2(f(x)), find g'(x). Leave answer in terms of e please!

OpenStudy (anonymous):

PLEASE help me out guys!

OpenStudy (anonymous):

\[e^{\ln(whatever)}=whatever\]

OpenStudy (anonymous):

and also \[e^{2\ln(\spadesuit)}=e^{\ln(\spadesuit^2)}=\spadesuit^2\]

OpenStudy (anonymous):

therefore \[\large e^{2\ln(1+\frac{1}{x})}=e^{\ln((1+\frac{1}{x})^2)}=(1+\frac{1}{x})^2\]

OpenStudy (anonymous):

and since \[(1+\frac{1}{x})^2=1+\frac{2}{x}+\frac{1}{x^2}\] taking the derivative should be okay